# SuperCoder: Assembly Program Superoptimization with Large Language Models

### v2

@@ 5.1 Evaluation of Different Models
Table 2: Comparison of LLMs on our assembly optimization benchmark. We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline. All open-source models are instruction-tuned.
@@
Table˜2 reports results across evaluated models. Most perform poorly on this task, with only a few demonstrating effectiveness as superoptimizers. Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates. Among all models, claude-opus-4 and claude-sonnet-4 perform best, achieving average speedups of 1.43× and 1.30×, respectively. Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler IRs. Among them, llm-compiler-13b achieves a notable 1.34× speedup, whereas the fine-tuned variants (-ftd) perform poorly, likely because they were adapted for different tasks such as disassembling x86-64 into LLVM IR. These results suggest that while superoptimization is inherently difficult, some LLMs can be effective superoptimizers.
@@ 5.2 Effectiveness of RL Training
Improvement. Table˜3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements. We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong test pass rate (61.4%) among models. After RL training with PPO, the fine-tuned model SuperCoder attains 95.0% correctness and improves average speedup from 1.10× to 1.46×. Its speedup percentiles are 1.17 ± 0.03× (25th), 1.35 ± 0.04× (50th), and 1.64 ± 0.08× (75th) respectively, outperforming the majority of evaluated models.
@@
Table 3: Performance of the base model and the models fine-tuned with RL. Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
5.3 Analysis of Learned Program Transformations
To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by our PPO-trained model SuperCoder. We categorize the learned transformations into four types: (1) code layout and instruction scheduling, which account for 98.5% of the changes and include basic block placement, instruction alignment, and reordering to improve cache behavior and hide latencies; (2) register allocation, representing 16% of the changes, where the model selects alternative registers; (3) control flow optimization (1.5%); and (4) instruction selection (1.0%). Since a single problem’s transformation may belong to multiple categories, the percentages do not sum to 100%. Unfortunately, there is currently no automated tool that can further reliably explain performance speedups at the assembly level. Automated analysis and explanation of performance differences in assembly code remains a challenging and unsolved problem.
The observed gains and our analysis suggest that LLMs can produce reasonable code transformations that meaningfully improve assembly performance. These results indicate that LLMs can uncover nuanced optimization opportunities beyond the reach of traditional compiler heuristics, suggesting that compilers themselves still have room for improvement.
5.4 Comparison with Simple Random Perturbations
To assess whether the improvements achieved by LLMs stem from systematic strategies learned during training rather than simple random variation in assembly programs, we compare against a baseline that applies simple random perturbations to assembly code. While such perturbations do not necessarily improve performance, running them multiple times and selecting the fastest version allows us to compute a speedup. These perturbations preserve correctness but can still influence performance, and we report speedup using the minimum observed runtime across 10 runs.
To randomly perturb the assembly, we compile the binaries with options that randomize both the stack and program load addresses. This alters code layout in memory, which can affect performance but not correctness. Speedup is measured as the ratio of the fastest run to the original runtime over 10 randomized runs. Repeating the experiment five times yields a speedup of 1.19 ± 0.01× (95% CI), significantly lower than 1.46 ± 0.12× achieved by SuperCoder. This indicates that SuperCoder systematically exploits performance opportunities rather than relying on simple random perturbations, particularly since it produces each result in a single shot without retries or iterative refinement.
@@ 6 Discussion
Besides the reward function presented in Section˜3.3, we also evaluate an alternative design. The original design assigns zero reward whenever any test fails. In contrast, the alternative assigns (i) a reward of 1-1 if the program fails to compile, (ii) a partial reward equal to the fraction of passed tests if only some tests succeed, and (iii) a reward of $1+\alpha\cdot\text{speedup}$ once all tests pass. Training the base model with this design yields an average speedup of 1.38×. Varying the scaling factor (5 or 10) has a negligible effect, and the result remains worse than the 1.46× achieved by SuperCoder with the original reward. This suggests that directly optimizing for the terminal speedup reward is more effective, since the alternative design allows models to avoid penalties by merely compiling or even gain positive reward from partially passing tests, which does not translate into actual speedup.
Direct Compilation from C.
We also examine a more challenging setting where gcc -O0 assembly is not provided. Instead, LLMs receive only the C source code and are asked to generate assembly directly. This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it. Similar degradation occurs for other models such as llm-compiler-13b and Claude models. These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
Case Study.
We illustrate in Appendix A.1 a representative example where an LLM discovers an optimization beyond the reach of a state-of-the-art compiler. The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit. The assembly produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches. In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation and reducing both instruction count and runtime overhead.
@@
A limitation of our approach is the lack of formal correctness guarantees. Although we validate generated programs with input–output test cases, testing is inherently incomplete and may miss corner cases even with high coverage. However, to our knowledge, no existing verifier can generally handle our assembly programs with loops. We hope this work motivates future research in the formal methods communities toward scalable verification techniques. The dataset we introduce could serve as a meaningful benchmark for such research.
Compared with traditional compiler optimization, reinforcement learning introduces substantial computational overhead during training. This cost, however, is incurred only once: once trained, the model can generate optimized assembly efficiently with low latency.
One promising direction for future work is to incorporate an interactive refinement loop, where the model iteratively improves its output based on feedback from errors or performance measurements. In addition, most superoptimization research has centered on x86-64, and our work follows this focus. Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability and impact. Such extensions would mainly require adapting the data collection and testing infrastructure, while keeping the core methodology unchanged.
@@ 7 Conclusion
We investigated whether LLMs can act as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. To support this study, we introduced the first large-scale benchmark of 8,072 assembly programs. Evaluating 23 models revealed the difficulty of the task, with most failing to achieve meaningful gains. By fine-tuning with reinforcement learning, our model SuperCoder improved both correctness and performance, reaching 95.0% test pass rate and an average speedup of 1.46× over gcc -O3. These results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly code, opening new opportunities for performance optimization beyond compiler heuristics.
LLM Usage
Large Language Models (LLMs) are the primary subject of study in this paper. In addition, we used LLMs as a general-purpose writing assistant to polish the presentation and improve readability.

### v3

@@ 5.1 Evaluation of Different Models
Table 2: Comparison of LLMs on our assembly optimization benchmark. We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline. Speedup is averaged across all test inputs, with each input evaluated over ten runs.
@@
Table 2 reports results across evaluated models. Most perform poorly on this task, with only a few demonstrating effectiveness as superoptimizers. Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates. Among all models, claude-opus-4 and claude-sonnet-4 perform best, achieving average speedups of 1.43× and 1.30×, respectively. Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler IRs. Among them, llm-compiler-13b achieves a notable 1.34× speedup, whereas the fine-tuned variants (-ftd) perform poorly, likely because they were adapted for different tasks such as disassembling x86-64 into LLVM IR. These results suggest that while superoptimization is inherently difficult, some LLMs can be effective superoptimizers.
@@ 5.2 Effectiveness of RL Training
Improvement. Table 3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements. We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong test pass rate (61.4%) among models. After RL training with PPO, the fine-tuned model SuperCoder attains 95.0% correctness and improves average speedup from 1.10× to 1.46×. Its speedup percentiles are 1.17 ± 0.03× (25th), 1.35 ± 0.04× (50th), and 1.64 ± 0.08× (75th) respectively, outperforming the majority of evaluated models.
@@
Table 3: Performance of the base model and the models trained with RL or supervised fine-tuning. Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
5.3 Results from Supervised Fine-Tuning and Inference-Time Methods
Best-of-N Sampling.
We evaluate best-of-N sampling for three models: claude-opus-4 (the strongest baseline in Table 2), Qwen2.5-Coder-7B (base), and SuperCoder (our PPO-trained model). Results are shown in Figure 3. Notably, the base model’s best-of-8 speedup is close to the PPO-trained model’s best-of-1 result, and the RL-trained model itself can still be improved with best-of-N sampling (i.e., from 1.46× in the single-sample setting to 1.93× with best-of-8 sampling).
Supervised Fine-Tuning.
We describe our supervised fine-tuning approach in Section 3.4. Table 2 reports results averaged over five runs with 95% confidence intervals. While supervised fine-tuning improves performance, RL achieves slightly stronger results. We believe that RL is a natural fit for the open-ended nature of superoptimization, as RL directly optimizes for correctness and speedup rather than imitating existing examples.
Iterative Refinement.
Figure 3 shows the results of iterative refinement, where the model receives feedback about compilation failures, test failures, or performance for self-reflection. All three models exhibit improvements as the number of refinement iterations increases, with the effect being most pronounced for our RL fine-tuned model.
Figure 2: Best-of-N sampling results.
Figure 3: Iterative refinement results.
5.4 Analysis of Program Transformations
To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by claude-opus-4 with best-of-8 sampling. We categorize the program transformations into nine types: 1) Loop restructuring covers changes to loop organization, including reordering, unrolling, or altered control flow; 2) Instruction selection captures the use of specialized CPU instructions (e.g., bsrq, popcnt, cmov) in place of longer generic instruction sequences; 3) Algorithmic simplification denotes replacing complex or custom logic with simpler algorithms or standard library routines such as memcmp, strcmp, or atoi; 4) Stack canary removal refers to eliminating stack protection checks and related security instrumentation; 5) Register allocation reflects differences in how registers are assigned, reused, or spilled to memory; 6) Branch elimination involves removing conditional branches by using conditional moves (cmov) or condition-setting instructions (setcc); 7) Address calculation optimization refers to simplifying memory address computations and offset arithmetic; 8) Function call optimization captures substituting heavyweight or checked function variants with simpler equivalents (e.g., __isoc99_scanf vs. scanf); and 9) Arithmetic optimization includes simplifying arithmetic operations such as replacing divisions or multiplications with shifts, or using increment/decrement instructions where applicable.
Appendix A.6 presents a detailed category-wise analysis, reporting the frequency of each transformation category among all transformed programs that pass all tests. For a given optimized version, the model may induce transformations spanning multiple optimization categories. Because performance optimization typically involves multiple interrelated decisions, it is difficult to automatically isolate which types of program transformations are responsible for observed performance improvements. More broadly, automated analysis and attribution of performance differences at the assembly level remains an open problem.
@@ 6 Discussion
Prompt Optimization Methods.
We experimented with few-shot in-context learning and found that adding more examples does not reliably improve performance and often degrades it, consistent with prior observations in code optimization (Shypula et al., 2024). We also evaluated GEPA (Agrawal et al., 2025), an evolutionary prompting framework that uses natural language reflection to derive optimization rules, but observed only minimal gains. Additional results are provided in Appendix A.4. Besides, we also experimented with prompts that include only the C source file, without providing the compiler-generated assembly. We find that omitting the assembly as a starting point leads to substantially worse results; these findings are discussed in Appendix A.5.
@@
Besides the reward function presented in Section 3.3, we also evaluate an alternative design. The original design assigns zero reward whenever any test fails. In contrast, the alternative assigns (i) a reward of 1-1 if the program fails to compile, (ii) a partial reward equal to the fraction of passed tests if only some tests succeed, and (iii) a reward of $1+\alpha\cdot\text{speedup}$ once all tests pass. Training the base model with this design yields an average speedup of 1.38×. Varying the scaling factor (5 or 10) has a negligible effect, and the result remains worse than the 1.46× achieved by SuperCoder with the original reward. This suggests that directly optimizing for the terminal speedup reward is more effective.
Case Studies.
We illustrate in Appendix A.1 two representative examples where an LLM discovers optimizations beyond the reach of a state-of-the-art compiler. The first example is about instruction selection. In the first example, the original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit. The assembly produced by gcc -O3 preserves this loop structure. In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation. The second example involves loop restructuring, along with function call optimization by replacing checked function variants with simpler equivalents.
@@
Our evaluation relies on test-based validation, which is common in prior work (Shypula et al., 2024; Du et al., 2024). While effective in practice, this approach would be more faithful with access to a formal equivalence checker for assembly programs. Existing state-of-the-art equivalence checkers for assembly (Schkufza et al., 2013) cannot handle programs with the rich control flow present in our dataset. To mitigate this limitation, our evaluation achieves 96.2% line coverage, substantially reducing the risk of undetected errors. Encouragingly, recent work has begun developing formal verification tools for GPU kernels (Dubey et al., 2025), partly motivated by advances in LLM-generated GPU code. We hope that our work further motivates the development of more general and scalable verification tools for assembly programs. In addition, most superoptimization research has centered on x86-64. Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability.
@@ 7 Conclusion
We investigated whether LLMs can act as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. To support this study, we introduced the first large-scale benchmark of 8,072 assembly programs. Evaluating 23 models revealed the difficulty of the task, with most failing to achieve meaningful gains. By fine-tuning with reinforcement learning, our model SuperCoder improved both correctness and performance, reaching 95.0% test pass rate and an average speedup of 1.46× over gcc -O3. We also show that Best-of-N sampling and iterative refinement can bring additional improvement. These results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly code, opening new opportunities for performance optimization beyond compiler heuristics.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ C Code
GCC -O3 Output
testl %eax, %eax
jle .L2
movl %eax, %r8d
jmp .L5
.p2align 4,,10
.p2align 3
.L4:
subl \$1, %r8d
je .L2
.L5:
movl %ecx, %eax
cltd
idivl %r8d
testl %edx, %edx
jne .L4
movl %r8d, %edx
leaq .LC1(%rip), %rsi
movl \$1, %edi
xorl %eax, %eax
call __printf_chk@PLT
.L2:
movq 8(%rsp), %rax
subq %fs:40, %rax
jne .L14
xorl %eax, %eax
addq \$24, %rsp
.cfi_remember_state
.cfi_def_cfa_offset 8
ret
Claude-Opus-4
testl %eax, %eax
jle .L2
.L3:
movl %ecx, %edx
movl %eax, %esi
movl %edx, %eax
cltd
idivl %esi
testl %edx, %edx
je .L4
decl %esi
movl %esi, %eax
jg .L3
.L2:
xorl %eax, %eax
addq \$24, %rsp
ret
.L4:
leaq .LC1(%rip), %rdi
xorl %eax, %eax
call printf@PLT
xorl %eax, %eax
addq \$24, %rsp
ret
.size main, .-main
Figure A2: Case study comparing the baseline assembly code snippet produced by gcc -O3 and the optimized assembly code snippet generated by Claude-Opus-4. Claude-Opus-4 eliminates the entry-path unconditional jump and alignment padding by fusing GCC’s two-block loop into a single, simpler control-flow structure, calls printf@PLT directly (a simpler function variant without security checks), and removes gcc’s stack-protector canary check.
@@ C Code:
A.4 Prompt Optimization Methods
Few-shot Prompting.
We evaluate 0-shot, 2-shot, and 4-shot prompting across three models. Table A2 shows that adding more in-context examples does not reliably improve performance and often degrades it. This observation is consistent with prior work (Shypula et al., 2024), which finds that few-shot examples can bias the model and lead to incorrect task understanding.
Model Shots Compile Pass (%) Test Pass (%) Avg. Speedup
claude-opus-4 0-shot 90.0 51.5 1.43×
claude-opus-4 2-shot 95.0 15.0 1.13×
claude-opus-4 4-shot 95.0 12.5 1.11×
SuperCoder (PPO) 0-shot 96.0 95.0 1.46×
SuperCoder (PPO) 2-shot 94.0 90.5 1.59×
SuperCoder (PPO) 4-shot 93.0 81.0 1.54×
Qwen2.5-Coder-7B (Base) 0-shot 77.9 61.4 1.10×
Qwen2.5-Coder-7B (Base) 2-shot 70.5 35.0 1.10×
Qwen2.5-Coder-7B (Base) 4-shot 80.5 30.5 1.06×
Table A2: Comparison of 0-shot, 2-shot, and 4-shot prompting across different models.
Figure A3: Categorization of the program transformations: Loop Restructuring (LR), Arithmetic Optimization (AO), Address Calculation (AC), Stack Canary Removal (SCR), Algorithmic Simplification (AS), Register Allocation (RA), Function Call Optimization (FC), Instruction Selection (IS), Branch Elimination (BE).
Prompt Evolution Framework.
We experimented with GEPA (Agrawal et al., 2025), an evolutionary prompting framework that uses natural language reflection to learn high-level optimization rules from trial and error. We used gpt-4o as the model under evaluation and gpt-5 as the reflection model. GEPA yielded only modest gains: compilation pass increased from 81.0% to 84.0% and test pass from 5.0% to 7.5%, while performance speedup remained essentially unchanged. We suspect this is because assembly optimization requires substantial domain knowledge that is difficult to capture by modifying the prompt alone.
A.5 Direct Compilation from C
We also examine a more challenging setting where gcc -O3 assembly is not provided. Instead, LLMs receive only the C source code and are asked to generate assembly directly. This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it. Similar degradation occurs for other models such as llm-compiler-13b and Claude models. These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
A.6 Categorization of Program Transformations
We categorize the differences between the code optimized by claude-opus-4 with best-of-8 sampling and the compiler-generated code into nine classes, each capturing a distinct kind of optimization or transformation: 1) Loop Restructuring (LR) covers changes to loop organization, including reordering, unrolling, or altered control flow; 2) Arithmetic Optimization (AO) includes simplifying arithmetic operations such as replacing divisions or multiplications with shifts, or using increment/decrement instructions where applicable; 3) Address Calculation (AC) optimization refers to simplifying memory address computations and offset arithmetic; 4) Stack Canary Removal (SCR) refers to eliminating stack protection checks and related security instrumentation; 5) Algorithmic Simplification (AS) denotes replacing complex or custom logic with simpler algorithms or standard library routines such as memcmp, strcmp, or atoi; 6) Register Allocation (RA) reflects differences in how registers are assigned, reused, or spilled to memory; 7) Function Call (FC) optimization captures substituting heavyweight or checked function variants with simpler equivalents (e.g., __isoc99_scanf vs. scanf); 8) Instruction Selection (IS) captures the use of specialized CPU instructions (e.g., bsrq, popcnt, cmov) in place of longer generic instruction sequences; 9) Branch Elimination involves removing conditional branches by using conditional moves (cmov) or condition-setting instructions (setcc). In  Figure A3, we plot the percentage of instances in our evaluation dataset that have each of these categories. Loop Restructuring (LR) is the most frequent category, appearing in 68.2% of instances, indicating that Claude often performs high-level changes to loop structure and control flow. Arithmetic Optimization (AO) and Address Calculation (AC) are also common, occurring in 45.0% and 30.5% of cases, respectively, while Stack Canary Removal (SCR) and Algorithmic Simplification (AS) appear in roughly a quarter of the dataset. Note that the total percentage exceeds 100% because an instance may have multiple transformation types.
A.7 Additional Evaluation on Randomly Sampled Programs
Our main dataset samples programs with larger speedup from -O0 to -O3 speedup. To evaluate how much dataset distribution affects our method, we conduct an additional experiment on 200 programs randomly sampled from CodeNet, and we ensure that this new set is fully disjoint from both our training and evaluation splits and not selected based on compiler speedup characteristics. Across these unbiased sampled programs, the conclusions remain largely unchanged: the PPO-trained model consistently produces correct and faster assembly code. This indicates that the improvements learned through reinforcement learning generalize beyond the specific distribution used for training and primary evaluation.
Model Compile Pass (%) Test Pass (%) Avg. Speedup
[…]
Table A3: Performance on 200 randomly sampled programs from CodeNet, disjoint from all training and evaluation splits. Unlike the main dataset, these programs are not selected for large gcc -O0 to gcc -O3 speedup. The PPO-trained model maintains strong correctness and performance gains.